10. A patient who is to undergo open heart surgery is to receive a continuous infusion of heparin during the procedure. The order is for 35 units of heparin per kg/hr. The patient weighs 172 lb. The heparin solution available is 20,000 units in 500 mL normal saline. What is the flow rate in mLs/hour for the ordered heparin solution?
Answer #1

The ordered dose is 35 unit of heparin per kg/hr.

The weight of patient is 172 lb. Therefore the weight of patient in kg can be calculated as

2.2 lb = 1 kg

172 lb = (1 x 172)/2.2 = 78.2 kg

So the patient weight is 78.2 kg.

Therefore the total unit of heparin ordered to that patient is

78.2 x 35 = 2737 units per hour.

Since the solution available is 20000 units in 500 mL, so we have to calculate the volume (X) that contains 2737 units.

=> 20000/500 = 2737/X

X = (500 x 2737)/20000

X = 68.42 mL

Therefore the rate of flow will be approximately 68.4 mL/hr